# Methods to Characterise Exoplanet Host Stars from Spectroscopy

### v1

@@ Abstract
A key to understand exoplanets is characterisation of their host stars. One of the most powerful tools to characterise stellar properties like effective temperature, surface gravity and metallicity, is spectroscopy based on observations of stellar atmospheres. This chapter describes the stellar parameters that can be derived from a spectrum with examples of well established methods and theoretical model atmospheres. Combined with photometry and parallax measurements, the outcome of the spectroscopic modelling can be used to derive stellar radii and masses.
@@ 1 Introduction
Exoplanets are intimately connected to their host stars through formation and evolution. In addition, detection and characterisation of exoplanets depend on detailed knowledge of their host stars since the current major detection techniques, transit photometry and the radial velocity (RV) method, detect planet sizes and masses relative to their host star (see Chapters by Deeg & Alonso and Wright in Volume 1 of the Handbook of Exoplanets). Uncertainties in a host star’s parameters propagate directly to the planets.
@@
Stellar modelling is normally based on two major techniques – photometry and spectroscopy. These methods are model-dependent in contrast to the direct measurements by interferometry, eclipsing binaries, and asteroseismology.
@@
For eclipsing binaries (e.g. Andersen 1991; Torres et al. 2010; Serenelli et al. 2021), and for the few large and nearby stars that enable interferometric measurements (e.g. Quirrenbach 2001; Eisenhauer et al. 2023), it is possible to derive the stellar radius with an accuracy of a few per cent. The timing of the duration of the eclipses of eclipsing binaries and their orbital velocities allows accurate estimates of their sizes. Similarly, stellar masses can be accurately determined for visual binaries from observed separations from the common centre-of-mass with Kepler’s third law that relates their masses with observed orbital period and separation. Asteroseismology can also be used to derive stellar mass, radius, and age with a high precision (see Chapter by Lundkvist, Huber, Aguirre & Chaplin in this volume of the Handbook of Exoplanets). Recently, the seismic surface gravity of the star has also been used to obtain the effective temperature and metallicity, in particular with APOGEE (Ahumada et al. 2020, 2022).
@@
These methods are, however, currently only possible to apply to a small subset of all stars. Spectroscopic measurements open a window to derive stellar parameters from a larger pool of stars than from direct methods. High-resolution spectroscopy is a powerful tool that provides a wealth of information; the effective temperature, surface gravity, chemical composition, and velocities. The stellar radius, and the luminosity, can then readily be derived from its spectral energy distribution (SED) via the spectroscopic parameters combined with photometry and parallax.
@@
The spectroscopic parameters also serve as a base to model characteristics that normally cannot be directly measured, like mass and age, with stellar evolution models and the complementary tool isochrones (e.g. Dotter 2016; Hidalgo et al. 2018). An isochrone is an evolutionary track for a population of stars with different masses on the Hertzsprung-Russell diagram with the same (iso) age (chrone). The mass and radius can also be obtained from the spectroscopic parameters and empirical calibration equations (e.g. Torres et al. 2010; Enoch et al. 2010; Southworth 2011) albeit often with a higher uncertainty.
@@
The downside of high-resolution spectroscopy is that it is expensive; the observations are time-consuming and also requires bright stars or a large collecting area. Modelling of stellar parameters has therefore traditionally been performed by photometry because a much larger number of stars can quickly be observed and analysed. The effective temperature of a star can for instance be derived from color-color diagrams which have been scaled to stars measured with direct methods (e.g. Bell and Gustafsson 1989; Alonso et al. 1996; Nordström et al. 2004). However, such models have in general higher uncertainties than based on spectroscopic measurements. In that respect they may only be reliable in a statistical sense and not for individual stars. Thus for exoplanet host stars (for which direct measurements are not applicable), high-resolution spectroscopy is preferred.
@@
This chapter begins with a short summary of a few basic requirements in order to obtain spectroscopic measurements of a star. It continues with an overview of the parameters that can be extracted from a high-resolution spectrum and a few examples of well established methods. Some advantages and caveats are highlighted. The chapter ends with a brief summary of how to combine spectroscopic parameters, photometric measurements, and stellar evolution models to obtain stellar radii and masses. Modelling of stellar ages are described by Christensen-Dalsgaard & Aguirre in a Chapter in this volume of the Handbook of Exoplanets.
@@ 2 Instruments and spectral resolution
Important properties of a spectrum is high-resolution, high signal-to-noise (S/N), and wavelength coverage. There are several different types of spectrometers and many textbooks have been written about this topic including advantages and challenges for different types of instruments (e.g. Gray 2008). The most successfull type in observations of exoplanets is echelle spectrographs. The main advantages is the high resolution combined with a wide wavelength coverage obtained in a single exposure. A description of high-precision cross-dispersed echelle spectrographs for exoplanet research (CORAVEL, ELODIE, CORALIE, SOPHIE and HARPS) is found in a Chapter by Fransesco Pepe in this volume of the Handbook of Exoplanets.
For the purpose of spectroscopic modelling of host stars, we want high-resolution in order to resolve the spectral lines. The spectral resolution at wavelength is
@@
Doppler broadening of spectral lines due to thermal and turbulent motions of absorbing species in the atmospheres produce line widths of km s-1 for late-type stars. This corresponds to a spectral resolution of 50 000. Low-mass, slowly rotating stars can have line widths of only km s-1 which require to resolve the spectral lines and disentangle blended lines.
In terms of high resolution, ultra-high precision, and long-term stability, the High Accuracy Radial Velocity Planet Searcher (HARPS; Mayor et al. 2003) and its decade younger sibling HARPS-North (Cosentino et al. 2012) mounted on the ESO 3.6 m telescope (La Silla observatory, Chile) and the Telescopio Nazionale Galileo (TNG) of Roque de los Muchachos Observatory (La Palma, Spain), respectively, have been the leading instruments in detecting exoplanets over the last two decades. Both HARPS instruments are fiber-fed cross-dispersed high-precision echelle spectrographs covering 380 nm to 690 nm. The spectroscopic resolution is approximately at visual wavelengths corresponding to a velocity resolution of 2.6 km s-1. The design is based on experience with the groundbreaking ELODIE and CORALIE instruments where the former was used to detect the first exoplanet 51 Peg b (Mayor and Queloz 1995). Both HARPS spectrographs can be considered as the “gold standard” when searching for exoplanets in RV data and are also used for characterisation of exoplanet host stars. When searching for exoplanets with the RV method, many measurements are collected, sometimes over a period over many years. The individual spectra can be co-added (after correcting for the periodic changes in radial velocity) in order to increase the S/N enabling spectroscopic characterisation of the host star.
In addition to the HARPS spectrographs, there are many other instruments used for exoplanet detection that can also be used to characterise host stars e.g. ESPRESSO (Pepe et al. 2021), FIES (Telting et al. 2014), HIRES (Vogt et al. 1994), CHIRON (Tokovinin et al. 2013), TULL (Tull et al. 1995) with different resolutions operating at different wavelengths.
@@ 3 Stellar properties from spectroscopy
Only a few parameters characterise a stellar atmosphere: the effective temperature (), the surface gravity (log ), the overall metal abundance ([M/H]), and atmospheric and rotational velocities.
The surface of a star is defined as the location where photons escape from the star which occurs at a characteristic optical depth of 2/3. This occurs within the photosphere, the innermost km of a star’s atmosphere which overlies the opaque interior. The photospheric temperature and density varies with depth and depends on the surface gravity as well as the abundances and opacity of the gases. The Sun’s photosphere has a temperature that varies between 4 400 K and 6 600 K with an effective temperature of 5 772 K, while the density is approximately kg m-3, increasing with depth into the Sun. Other stars may have hotter or cooler photospheres. Spectral absorption lines originate from different depths and opacities within the photosphere. Weak and optically thin spectral lines, and the line wings of optically thick lines, originate essentially in the same layer as the continuum. In contrast, the cores of optically thick saturated lines develop in higher layers. An example is the core of the hydrogen line at 6562.81 Å which originates in the hotter and less dense chromosphere which lies on top of the photosphere, where the assumption of local thermal equilibrium (LTE) is no longer valid.
@@
Instead of choosing a particular depth to define a star’s surface temperature, the effective temperature is defined in terms of flux. The effective temperature is defined via the Stefan-Bolztmann law in terms of the total power per unit area, radiated by the star (e.g. Gray 2008)
@@
where the left integral is the total radiative flux from the star received at the top of the Earth’s atmosphere (bolometric flux), is the stellar radius, is the distance, and is the angular diameter of the star. The effective temperature can hence be derived by measuring the angular size of the star from interferometry and the received flux at Earth over a wide spectral range. By combining the angular size with distance from parallax measurements, the linear radius of the star can be inferred for a wide range of spectral types nearly model-independent except for the dependence on the adopted limb-darkening coefficients and bolometric correction. Alternatively, if the radius of the star is known from e.g. eclipsing binaries and the distance from parallax measurments, this will give the angular size which then can be used to compute . Since the target stars have to be nearby to measure their angular sizes, interstellar absorption can be neglected. The infrared flux method (IRFM Blackwell and Shallis 1977; Blackwell et al. 1980; Bell and Gustafsson 1989; Blackwell et al. 1990; Casagrande et al. 2006, 2010) is based on observations of the angular size of the star and the measured infrared flux at the top of the Earth’s atmosphere. The bolometric flux is derived taking into account the bolometric correction which gives the effective temperature.
@@
The surface gravity of a star is defined by (e.g. Gray 2008)
@@
A stars chemical composition is an outcome of the nucleosynthesis by previous generations of stars. This is important when reconstructing star and planet formation history in our Galaxy. There is a large variation of metallicity, i.e. the abundance of all elements heavier than helium denoted with [M/H], in the Milky Way up to about twice the solar value to hundreds of thousands of times lower than the solar value (e.g. Christlieb et al. 2004; Li et al. 2022; Nepal et al. 2024).
The chemical composition of a star is commonly fixed to the overall metallicity of a star relative to the Sun. The Grevesse et al. (2007), Asplund et al. (2009) and Lodders (2003) abundance scales for the Sun are currently the most adopted. However, individual abundances of a star may not follow solar composition and may require modelling of individual elemental abundances. Abundances are generally measured on a logarithmic scale normalised to the Sun where zero equals the Sun’s metallicity. In the case of iron we have [Fe/H]⋆ = log(Fe/H)⋆ - log(Fe/H)⊙. For example, an iron abundance of [Fe/H] means that the abundance is relative to the Sun. Since iron is by far the most abundant species in a stellar atmosphere after hydrogen and helium, measurements of iron has become a proxy for the metallicity.
Stellar abundances is also important when modelling exoplanet interiors in particular rocky super-Earths without significant gaseous envelopes. The degeneracy of interior composition inferred from radius and mass measurements can for this type of planet be reduced assuming an interior structure with a differentiated iron core and a rocky mantle. In these cases, the host star abundances is often used as a proxy of the primary planet-building elements Fe, Mg, and Si, which are expected to be reflected in the planet composition, planet interior, and core mass fraction (Dorn et al. 2015; Acuña et al. 2023).
@@ 4 Methods for spectroscopic modelling
There are several ways to model a spectrum which can be divided into two main groups. The first is based on spectral synthesis. Here observations are fitted to a synthetic spectrum of stellar atmosphere models by comparison of line profiles. The second is a line-by-line analysis based on measured strengths of observed spectral lines, their equivalent widths (EWs). Detailed description of the physics can be found in many textbooks e.g. Gray (2008). Spectroscopic observations can also be compared to a library of spectra of well-characterised stars via for example interferometric measurements or spectroscopic binaries.
One major problem in spectroscopic analysis is to accurately determine the continuum which can introduce large errors. This is particularly difficult for poor spectra with low spectral resolution or low S/N. It also depends on the spectral type of the star and the wavelength region. The number of spectral lines increases towards shorter wavelengths for all types of stars. In addition, late-type stars have a much higher density of spectral lines than early-type stars arising from both atoms and molecules leading to blending and confusion of the continuum location. The higher temperatures of early-type stars ionize a large fraction of their atoms, leading to significantly fewer spectral features than low-mass stars. Differences in rotational velocities also affect the spectral line density. In contrast to late-type stars, the early types have very high rotational velocities which leads to very broad spectral lines that smear out spectral features. Thus solar-type stars (FGK) are the easiest stars to model, while high- and low-mass stars often entails a significantly higher degree of difficulty in the modelling. M-dwarfs have in addition generally a much longer period of high stellar activity than FGK stars, exacerbating the problems (e.g. Mignon et al. 2023). This is unfortunate since M-dwarfs are popular exoplanet host stars due to their small masses and sizes which increase the exoplanet signals.
Care must be taken when selecting which spectral lines to model. A large set of narrow, non-blended spectral lines are preferred (unless modelling pressure broadened line wings, see below). If a spectral line becomes optically thick, the abundance of a species stops growing linearly with absorption depth. The characteristics of an optically thick line is a saturated line centre which flattens the bottom and broaden the line wings. Not all optically thin spectral lines may, however, be useful since a large number comes with poorly determined atomic parameters which are needed to compute synthetic spectra. This can be circumvented for solar-type stars if adopting new atomic parameters after comparing the lines from observations of the Sun.

### v2

# Methods to Characterize Exoplanet Host Stars from Spectroscopy

@@ Abstract
A key to understand exoplanets is characterization of their host stars. One of the most powerful tools to characterize stellar properties like effective temperature, surface gravity and metallicity, is spectroscopy based on observations of stellar atmospheres. This chapter describes the stellar parameters that can be derived from a spectrum with examples of well established methods and theoretical model atmospheres. Combined with photometry and parallax measurements, the outcome of the spectroscopic modelling can be used to derive stellar radii and masses.
@@ 1 Introduction
Exoplanets are intimately connected to their host stars through formation and evolution. In addition, detection and characterization of exoplanets depend on detailed knowledge of their host stars since the current major detection techniques, transit photometry, and the radial velocity (RV) method detect planet sizes and masses relative to their host star as described in Radial Velocities as an Exoplanet Discovery Method (Jason T. Wright) and Transit Photometry as an Exoplanet Discovery Method (Hans Deeg and Roi Alonso). Uncertainties in a host star’s parameters propagate directly to the planets.
@@
Stellar modelling is normally based on two major techniques – photometry and spectroscopy. These methods are model dependent in contrast to the direct measurements by interferometry, eclipsing binaries, and asteroseismology.
@@
For eclipsing binaries (e.g., Andersen 1991; Torres et al. 2010; Serenelli et al. 2021), and for the few large and nearby stars that enable interferometric measurements (e.g., Quirrenbach 2001; Eisenhauer et al. 2023), it is possible to derive the stellar radius with an accuracy of a few percent. The timing of the duration of the eclipses of eclipsing binaries and their orbital velocities allows accurate estimates of their sizes. Similarly, stellar masses can be accurately determined for visual binaries from observed separations from the common center of mass with Kepler’s third law that relates their masses with observed orbital period and separation. Asteroseismology can also be used to derive stellar mass, radius, and age with a high precision as described in detail in Characterizing Host Stars Using Asteroseismology by Mia Sloth Lundkvist, Daniel Huber, Víctor Silva Aguirre, and William J. Chaplin. Recently, the seismic surface gravity of the star has also been used to obtain the effective temperature and metallicity, in particular with APOGEE (Ahumada et al. 2020, 2022).
@@
These methods are, however, currently only possible to apply to a small subset of all stars. Spectroscopic measurements open a window to derive stellar parameters from a larger pool of stars than from direct methods. High-resolution spectroscopy is a powerful tool that provides a wealth of information: the effective temperature, surface gravity, chemical composition, and velocities. The stellar radius, and the luminosity, can then readily be derived from its spectral energy distribution (SED) via the spectroscopic parameters combined with photometry and parallax.
@@
The spectroscopic parameters also serve as a base to model characteristics that normally cannot be directly measured, like mass and age, with stellar evolution models and the complementary tool isochrones (e.g., Dotter 2016; Hidalgo et al. 2018). An isochrone is an evolutionary track for a population of stars with different masses on the Hertzsprung-Russell diagram with the same (iso) age (chrone). The mass and radius can also be obtained from the spectroscopic parameters and empirical calibration equations (e.g., Torres et al. 2010; Enoch et al. 2010; Southworth 2011) albeit often with a higher uncertainty.
@@
The downside of high-resolution spectroscopy is that it is expensive; the observations are time-consuming and also require bright stars or a large collecting area. Modelling of stellar parameters has therefore traditionally been performed by photometry because a much larger number of stars can quickly be observed and analyzed. The effective temperature of a star can for instance be derived from color-color diagrams which have been scaled to stars measured with direct methods (e.g., Bell and Gustafsson 1989; Alonso et al. 1996; Nordström et al. 2004). However, such models have in general higher uncertainties than based on spectroscopic measurements. In that respect they may only be reliable in a statistical sense and not for individual stars. Thus for exoplanet host stars (for which direct measurements are not applicable), high-resolution spectroscopy is preferred.
@@
This chapter begins with a short summary of a few basic requirements in order to obtain spectroscopic measurements of a star. It continues with an overview of the parameters that can be extracted from a high-resolution spectrum and a few examples of well-established methods. Some advantages and caveats are highlighted. The chapter ends with a brief summary of how to combine spectroscopic parameters, photometric measurements, and stellar evolution models to obtain stellar radii and masses. Modelling of stellar ages is described in Ages for Exoplanet Host Stars by Christensen-Dalsgaard and Aguirre.
@@ 2 Instruments and spectral resolution
Important properties of a spectrum are high-resolution, high signal to noise (S/N), and wavelength coverage. There are several different types of spectrometers, and many textbooks have been written about this topic including advantages and challenges for different types of instruments (e.g., Gray 2008). The most successful type in observations of exoplanets is echelle spectrographs. The main advantage is the high resolution combined with a wide wavelength coverage obtained in a single exposure. Francesco Pepe describes the high-precision cross-dispersed echelle spectrographs for exoplanet research CORAVEL, ELODIE, CORALIE, SOPHIE, and HARPS in High-Precision Spectrographs for Exoplanet Research.
For the purpose of spectroscopic modelling of host stars, we want high resolution in order to resolve the spectral lines. The spectral resolution at wavelength is
@@
Doppler broadening of spectral lines due to thermal and turbulent motions of absorbing species in the atmospheres produces line widths of km s-1 for late-type stars. This corresponds to a spectral resolution of 50,000. Low-mass, slowly rotating stars can have line widths of only km s-1, which require to resolve the spectral lines and disentangle blended lines.
In terms of high resolution, ultra-high precision, and long-term stability, the High-Accuracy Radial Velocity Planet Searcher (HARPS; Mayor et al. 2003) and its decade younger sibling HARPS-North (Cosentino et al. 2012) mounted on the ESO 3.6 m telescope (La Silla observatory, Chile) and the Telescopio Nazionale Galileo (TNG) of Roque de los Muchachos Observatory (La Palma, Spain), respectively, have been the leading instruments in detecting exoplanets over the last two decades. Both HARPS instruments are fiber-fed cross-dispersed high-precision echelle spectrographs covering 380 to 690 nm. The spectroscopic resolution is approximately at visual wavelengths corresponding to a velocity resolution of 2.6 km s-1. The design is based on experience with the groundbreaking ELODIE and CORALIE instruments where the former was used to detect the first exoplanet 51 Peg b (Mayor and Queloz 1995). Both HARPS spectrographs can be considered as the “gold standard” when searching for exoplanets in RV data and are also used for characterization of exoplanet host stars. When searching for exoplanets with the RV method, many measurements are collected, sometimes over a period over many years. The individual spectra can be co-added (after correcting for the periodic changes in radial velocity) in order to increase the S/N enabling spectroscopic characterization of the host star.
In addition to the HARPS spectrographs, there are many other instruments used for exoplanet detection that can also be used to characterize host stars e.g. ESPRESSO (Pepe et al. 2021), FIES (Telting et al. 2014), HIRES (Vogt et al. 1994), CHIRON (Tokovinin et al. 2013), TULL (Tull et al. 1995) with different resolutions operating at different wavelengths.
@@ 3 Stellar Properties from Spectroscopy
Only a few parameters characterize a stellar atmosphere: the effective temperature (), the surface gravity (log ), the overall metal abundance ([M/H]), and atmospheric and rotational velocities.
The surface of a star is defined as the location where photons escape from the star which occurs at a characteristic optical depth of 2/3. This occurs within the photosphere, the innermost km of a star’s atmosphere which overlies the opaque interior. The photospheric temperature and density varies with depth and depend on the surface gravity as well as the abundances and opacity of the gases. The Sun’s photosphere has a temperature that varies between 4400 and 6600 K with an effective temperature of 5772 K, while the density is approximately kg m-3, increasing with depth into the Sun. Other stars may have hotter or cooler photospheres. Spectral absorption lines originate from different depths and opacities within the photosphere. Weak and optically thin spectral lines, and the line wings of optically thick lines, originate essentially in the same layer as the continuum. In contrast, the cores of optically thick saturated lines develop in higher layers. An example is the core of the hydrogen line at 6562.81 Å which originates in the hotter and less dense chromosphere which lies on top of the photosphere, where the assumption of local thermal equilibrium (LTE) is no longer valid.
@@
Instead of choosing a particular depth to define a star’s surface temperature, the effective temperature is defined in terms of flux. The effective temperature is defined via the Stefan-Boltzmann law in terms of the total power per unit area, radiated by the star (e.g., Gray 2008)
@@
where the left integral is the total radiative flux from the star received at the top of the Earth’s atmosphere (bolometric flux), is the stellar radius, is the distance, and is the angular diameter of the star. The effective temperature can hence be derived by measuring the angular size of the star from interferometry and the received flux at Earth over a wide spectral range. By combining the angular size with distance from parallax measurements, the linear radius of the star can be inferred for a wide range of spectral types nearly model independent except for the dependence on the adopted limb-darkening coefficients and bolometric correction. Alternatively, if the radius of the star is known from e.g., eclipsing binaries and the distance from parallax measurements, this will give the angular size which then can be used to compute . Since the target stars have to be nearby to measure their angular sizes, interstellar absorption can be neglected. The infrared flux method (IRFM; Blackwell and Shallis 1977; Blackwell et al. 1980, 1990; Bell and Gustafsson 1989; Casagrande et al. 2006, 2010) is based on observations of the angular size of the star and the measured infrared flux at the top of the Earth’s atmosphere. The bolometric flux is derived taking into account the bolometric correction which gives the effective temperature.
@@
The surface gravity of a star is defined by (e.g., Gray 2008)
@@
A stars chemical composition is an outcome of the nucleosynthesis by previous generations of stars. This is important when reconstructing star and planet formation history in our Galaxy. There is a large variation of metallicity, i.e., the abundance of all elements heavier than helium denoted with [M/H], in the Milky Way up to about twice the solar value to hundreds of thousands of times lower than the solar value (e.g., Christlieb et al. 2004; Li et al. 2022; Nepal et al. 2024).
The chemical composition of a star is commonly fixed to the overall metallicity of a star relative to the Sun. The Grevesse et al. (2007), Asplund et al. (2009) and Lodders (2003) abundance scales for the Sun are currently the most adopted. However, individual abundances of a star may not follow solar composition and may require modelling of individual elemental abundances. Abundances are generally measured on a logarithmic scale normalized to the Sun where zero equals the Sun’s metallicity. In the case of iron we have [Fe/H]⋆ = log(Fe/H)⋆ - log(Fe/H)⊙. For example, an iron abundance of [Fe/H] means that the abundance is relative to the Sun. Since iron is by far the most abundant species in a stellar atmosphere after hydrogen and helium, measurements of iron have become a proxy for the metallicity.
Stellar abundances are also important when modelling exoplanet interiors in particular rocky super-Earths without significant gaseous envelopes. The degeneracy of interior composition inferred from radius and mass measurements can for this type of planet be reduced assuming an interior structure with a differentiated iron core and a rocky mantle. In these cases, the host star abundances are often used as a proxy of the primary planet-building elements Fe, Mg, and Si, which are expected to be reflected in the planet composition, planet interior, and core mass fraction (Dorn et al. 2015; Acuña et al. 2023).
@@ 4 Methods for Spectroscopic Modelling
There are several ways to model a spectrum which can be divided into two main groups. The first is based on spectral synthesis. Here observations are fitted to a synthetic spectrum of stellar atmosphere models by comparison of line profiles. The second is a line-by-line analysis based on measured strengths of observed spectral lines and their equivalent widths (EWs). Detailed description of the physics can be found in many textbooks e.g., Gray (2008). Spectroscopic observations can also be compared to a library of spectra of well-characterized stars via for example interferometric measurements or spectroscopic binaries.
One major problem in spectroscopic analysis is to accurately determine the continuum which can introduce large errors. This is particularly difficult for poor spectra with low spectral resolution or low S/N. It also depends on the spectral type of the star and the wavelength region. The number of spectral lines increases toward shorter wavelengths for all types of stars. In addition, late-type stars have a much higher density of spectral lines than early-type stars arising from both atoms and molecules leading to blending and confusion of the continuum location. The higher temperatures of early-type stars ionize a large fraction of their atoms, leading to significantly fewer spectral features than low-mass stars. Differences in rotational velocities also affect the spectral line density. In contrast to late-type stars, the early types have very high rotational velocities which leads to very broad spectral lines that smear out spectral features. Thus solar-type stars (FGK) are the easiest stars to model, while high- and low-mass stars often entail a significantly higher degree of difficulty in the modelling. M-dwarfs have in addition generally a much longer period of high stellar activity than FGK stars, exacerbating the problems (e.g., Mignon et al. 2023). This is unfortunate since M-dwarfs are popular exoplanet host stars due to their small masses and sizes which increase the exoplanet signals.
Care must be taken when selecting which spectral lines to model. A large set of narrow, non-blended spectral lines are preferred (unless modelling pressure broadened line wings, see below). If a spectral line becomes optically thick, the abundance of a species stops growing linearly with absorption depth. The characteristics of an optically thick line is a saturated line center which flattens the bottom and broaden the line wings. Not all optically thin spectral lines may, however, be useful since a large number comes with poorly determined atomic parameters which are needed to compute synthetic spectra. This can be circumvented for solar-type stars if adopting new atomic parameters after comparing the lines from observations of the Sun.